The acceleration of a particle is increasing linearly with time t as bt.the particle starts from the origin with an initial velocity V0.the distance travelled by the particle in time t will be
 

Grade:10

1 Answers

Arun
5 years ago
 

d2s/dt2 =bt

intigrate

V= ds/dt = bt2/2 +c

at t=0 V=U

so U =0 +c

ds/dt = bt2/2 +U

furthet intigrate

s= b/6 t3 +Ut +d

at t=0 s=0

so d=0

S =b/6 t3 +Ut

 

Regards

arun